OpenStudy (anonymous):

how do you find the limits of a double integral if they give you y = x^2 and y = 2x ?

OpenStudy (sirm3d):

solve the system of equations \(y=x^2\) and \(y=2x\) these values will be your \(x_L\) and \(x_R\)  \(y_B=x^2, y_T=2x\) \[\Large \iint_R f(x,y) \mathrm dA=\int_{x_L}^{x_R} \int_{y_B}^{y_T} f(x,y) \mathrm dy \mathrm dx\]

OpenStudy (anonymous):

would the limits of x be 0,2 and y would be from x^2 to 2x ??